# HPE-CogVLM: New Head Pose Grounding Task Exploration on Vision Language Model

### v1

@@ Abstract
Head pose estimation (HPE) task requires a sophisticated understanding of 3D spatial relationships and precise numerical output of yaw, pitch, and roll Euler angles. Previous HPE studies are mainly based on Non-large language models (Non-LLMs), which rely on close-up human heads cropped from the full image as inputs and lack robustness in real-world scenario. In this paper, we present a novel framework to enhance the HPE prediction task by leveraging the visual grounding capability of CogVLM. CogVLM is a vision language model (VLM) with grounding capability of predicting object bounding boxes (BBoxes), which enables HPE training and prediction using full image information input. To integrate the HPE task into the VLM, we first cop with the catastrophic forgetting problem in large language models (LLMs) by investigating the rehearsal ratio in the data rehearsal method. Then, we propose and validate a LoRA layer-based model merging method, which keeps the integrity of parameters, to enhance the HPE performance in the framework. The results show our HPE-CogVLM achieves a 31.5% reduction in Mean Absolute Error for HPE prediction over the current Non-LLM based state-of-the-art in cross-dataset evaluation. Furthermore, we compare our LoRA layer-based model merging method with LoRA fine-tuning only and other merging methods in CogVLM. The results demonstrate our framework outperforms them in all HPE metrics.
**footnotetext: Corresponding authors.
1 Introduction
@@
Nowadays, the head pose estimation (HPE) technique is applicable in various fields such as attention estimation fischer2018rt ; kellnhofer2019gaze360 ; cheng2018appearance , face recognition 7780892 ; 8255649 ; 5959981 ; 7299081 , customer behavior analysis 6655785 ; wu2016head , driver assistance systems 5443483 ; vora2018driver ; hu2021temporal ; hu2020robust and human-robot interaction app11125366 . This task involves predicting the Euler angles (yaw, pitch, and roll) of human heads from images or videos. Recent research efforts on some Non-Large Language Models (Non-LLMs) like 6DRepNet Hempel_2022 , HopeNet DBLP:journals/corr/abs-1710-00925 and WHENet zhou2020whenet have made significant advancements in HPE.
@@
Despite the recent surge of interest in HPE, the application of this technique still faces several challenges in real-world scenario. The Non-LLMs typically rely on narrowly focused datasets such as 300W-LP Zhu_2019 for training, and validate on similarly constrained datasets like AFLW2000 Zhu_2019 and BIWI 6553713 . These datasets mainly feature close-up images of heads, mostly displaying frontal faces with yaw angles from $-99^{\circ}$ to $99^{\circ}$, instead of covering the entire range of head poses ranged from $-180^{\circ}$ to $180^{\circ}$. Additionally, the frequent use of close-up images in these datasets not only leads to uniform backgrounds, but also reduces the variability in the input data. The uniformity from datasets results in a lack of robustness in diverse real-world environments. The DirectMHP zhou2023directmhp model achieves the HPE prediction in a one-shot manner trained on the full-range HPE datasets, like Agora patel2021agora and CMU joo2016panoptic , but this model struggles to balance the head bounding box (BBox) detection and HPE task performance. Consequently, the model’s effectiveness remains uncertain in real-world settings.
@@
Large Language Models (LLMs) significantly improve our lives by offering sophisticated assistance in various tasks. Recently, Vision Language Models (VLMs) have gained prominence due to their proficiency in interpreting and processing information from images and videos openai2024gpt4 ; geminiteam2024gemini ; liu2023visual ; wang2024cogvlm . Through integrating visual capabilities to LLMs, the VLMs can achieve more complex tasks than traditional LLMs, e.g., visual question answering openai2024gpt4 ; agrawal2016vqa ; liu2023visual and visual grounding wang2024cogvlm ; yu2016modeling . The visual grounding ability of CogVLM wang2024cogvlm exhibits strong adaptability to diverse environments, providing an opportunity to enhance the robustness of tasks that traditional CNN-based methods struggle to solve. In this paper, we aim to improve HPE task functionality with the grounding CogVLM. The grounding CogVLM’s capabilities include caption grounding, referring expression generation, referring expression comprehension and grounded visual question answering wang2024cogvlm . All of these functionalities involve in the description of object localization in the BBox format of [[$x_{0}$, $y_{0}$, $x_{1}$, $y_{1}$]] as shown in Figure 1 (a). This BBox prediction capability provides a foundational skill for learning the new HPE task in this paper. By leveraging this capability in our designed prompts, we enable CogVLM to learn HPE from the entire images instead of the cropped head images used in Non-LLMs, which significantly helps the model avoid over-fitting to a limited background.
@@
Integrating the HPE task into the grounding CogVLM not only opens up new opportunities for exploration, but also presents several challenges. First, VLM tasks such as image description, visual reasoning, and visual perception usually contain answering questions with natural language responses. In contrast, our HPE task requires the VLM to produce precise numerical Euler angles. Although the grounding CogVLM can predict BBoxes, indicating its ability to produce numerical responses, the HPE task is significantly more complicated. HPE requires predicting the human head’s orientation in terms of yaw, pitch, and roll angles, which involves interpreting 3D orientation from 2D images, introducing additional dimensions of depth and angular perspective not required in the basic BBox detection task. Therefore, it raises the challenge of whether the grounding model can provide HPE answers with much higher accuracy. Secondly, catastrophic forgetting scialom2022finetuned ; huang2024mitigating ; luo2024empirical poses a significant challenge in fine-tuning LLMs. The catastrophic forgetting problem is a phenomenon that LLMs tend to forget previously learned information when acquiring new data. Currently, there is a lack of research addressing the catastrophic forgetting problem in complex grounding tasks. Lastly, the original grounding CogVLM only involves in outputting responses with a blend of natural languages and BBoxes in [[$x_{0}$, $y_{0}$, $x_{1}$, $y_{1}$]] format. In this paper, we introduce a new format {$yaw\_angle$, $pitch\_angle$, $roll\_angle$} for answering HPE prompts as shown in Figure 1 (b). This enriches the knowledge of the original grounding CogVLM, meanwhile increasing the complexity of output formats. Empirically, we observe the LoRA hu2021lora fine-tuning and model merging methods often generate invalid blend outputs like [[$x_{0}$, $y_{0}$, $yaw\_angle$}, which is referred as invalid answers in this paper. More invalid answers are detailed in Appendix A.1 Table 5.
@@
Figure 1: Figure (a) shows an example of CogVLM Grounding Capability, which demonstrates the original grounding CogVLM’s ability to identify objects based on prompts, a foundational skill useful for HPE task. Figure (b) displays a visualization of head orientation predicted by our HPE-CogVLM from the CMU Panoptic dataset, using Euler angles. The head pose labels are depicted with pitch (red axis), roll (green axis), and yaw (blue axis) angles, each indicated in their respective directions.
@@
In this paper, for addressing the catastrophic forgetting problem in grounding tasks, we evaluate and improve the data rehearsal methods scialom2022finetuned ; huang2024mitigating used in non-grounding VLMs to overcome the catastrophic forgetting problem. Here, the rehearsal ratio represents the percentage of images randomly selected from earlier training phases that are reintegrated during the training of new tasks, relative to the total number of earlier training images scialom2022finetuned ; huang2024mitigating . The results show that the visual grounding task, which demands multiple accurate numerical outputs, requires a significantly larger rehearsal ratio than non-grounding VLMs. We propose and validate a layer-based model merging method to enhance the HPE task’s performance. Utilizing this merging approach, our model shows remarkable robustness, achieving a 31.5% reduction in Mean Absolute Error (MAE) of Euler angles, compared to the Non-LLMs state-of-the-art (SOTA) in cross-dataset evaluations. Furthermore, we compare our layer-based model merging method with LoRA fine-tuned and merged models in CogVLM. Our approach consistently shows superior performance in both of MAE and invalid answer ratio reduction. Our contributions can be concluded as following:
@@
  • Our work pioneers the improvement of HPE tasks through leveraging the visual grounding capability of CogVLM, showing the VLM model’s ability at managing complex 3D spatial perception while keeping the existing object localization knowledge.
  • We first explore the catastrophic forgetting problem and the invalid answer problem in the complex VLM grounding tasks.
@@
  • We propose a novel layer-based model merging method that adopts a “winner takes all” strategy, which significantly outperforms the Non-LLMs SOTA and VLM-based model in MAE and invalid answer ratio reduction. This demonstrates our method is outstanding robustness and effectiveness in the HPE task, and holds potential for broader application in various grounding tasks. The paper is under review and codes will be released.
2 Related Work
Head Pose Estimation. Traditional approaches for HPE include landmark-based and landmark-free methods. Since full-range HPE often involves head orientations where facial features are not visible, we focus solely on landmark-free approach. Under this approach, several models divide continuous rotation variables into discrete bins for classification purposes DBLP:journals/corr/abs-1710-00925 ; zhou2020whenet ; 8444061 ; huang2020improving ; zhang2020fdn . Besides those, FSA-Net yang2019fsa employs a stage-wise regression and feature aggregation scheme to predict Euler angles. 6DRepNet Hempel_2022 and TriNet cao2020vectorbased estimate the rotation matrix rather than Euler angles. These Non-LLM methods pose significant robustness issues in real-life scenarios.
Grounding Vision Language Models. Some VLMs with grounding capabilities can provide accurate BBox information in the format of [[$x_{0}$, $y_{0}$, $x_{1}$, $y_{1}$]] wang2024cogvlm ; bai2023qwenvl ; chen2023shikra ; you2023ferret . In this paper, we focus a unique grounding task: HPE prediction that requires a understanding of 3D spatial relationships to produce accurate Euler angles in the format of {$yaw\_angle$, $pitch\_angle$, $roll\_angle$}. Regular VLMs are not typically designed to handle such queries, and currently there is a lack of research investigating the effectiveness of VLMs in HPE task. The research of CLIP-Gaze yin2024clip , learns the gaze estimation task by using CLIP radford2021learning model. However, CLIP-Gaze limits its application solely to gaze estimation and does not address catastrophic forgetting, resulting in a loss of CLIP’s multi-task capabilities.
Catastrophic Forgetting Problem. Catastrophic forgetting has been a significant issue that limits the effectiveness of LLMs, as they tend to forget previously knowledge when learning new knowledge. Kirkpatrick et al. Kirkpatrick_2017 and Li et al. li-etal-2022-overcoming control the extent of parameter updates to prevent the forgetting problem of previously learned tasks, which needs to carefully tune for optimal performance. Xu et al. xu2018reinforced and Huang et al. huang2021continual separate parameters dedicated to individual tasks, however, this approach introduces extra parameters. The rehearsal method scialom2022finetuned ; huang2024mitigating ; luo2024empirical ; zhang-etal-2023-citb ; mok-etal-2023-large is the most widely used method to mitigate catastrophic forgetting which reuses a small portion of old task datasets into the new task fine-tuning process. The catastrophic forgetting problem is not well investigated for complex grounding task in previous literature. In this paper, we evaluate and improve rehearsal method for our grounding tasks.
Model Merging in LLMs. There has been extensive exploration of methods based on model merging to enhance the performance of LLMs. This method merges multiple LLMs with specialized capabilities, into a single LLM that can address tasks across various domains. The typical merging methods ilharco2023editing ; wortsman2022model ; jang2024model ; davari2023model ; yadav2024ties ; yu2023language ; akiba2024evolutionary usually apply rules or algorithms to trim or merge the parameters of LLMs. For example, task arithmetic ilharco2023editing defines arithmetic rules to incorporating new capability or deleting undesired capability. Evolutionary model merging method akiba2024evolutionary has demonstrated evolutionary algorithms’ effectiveness in LLMs. However, traditional merging methods usually produce blended invalid answers as our tasks requires higher complexity of output formats.
3 HPE-CogVLM Framework
@@
The proposed framework of HPE-CogVLM as shown in Figure 2 is structured through a multi-stage process. It is specially designed to enhance the model’s capability in understanding and processing complex tasks associated with HPE, while retaining its original BBox prediction capabilities. The fine-tuning process at each stage follows the CogVLM’s fine-tuning scripts [^1] https://github.com/THUDM/CogVLM, which implement LoRA hu2021lora across transformer blocks, including the query, key, value of attention layers and dense layers. Subsequently, the LoRA matrices of each layer are accumulated into the corresponding layer in original model. Below is a detailed description of each stage in the framework:
@@
Stage 1: Pre-training of the Original Grounding CogVLM on Weak Label Data

### v2

# HPE-CogVLM: Advancing Vision Language Models with a Head Pose Grounding Task

@@ Abstract
Head pose estimation (HPE) requires a sophisticated understanding of 3D spatial relationships to generate precise yaw, pitch, and roll angles. Previous HPE models, primarily CNN-based, rely on cropped close-up human head images as inputs and often lack robustness in real-world scenario. Vision Language Models (VLMs) can analyze entire images while focusing on specific objects through their attention mechanisms. In this paper, we propose a novel framework to improve the HPE accuracy by leveraging the object detection grounding capability of a VLM, referred to as CogVLM. We empirically find that directly LoRA fine-tuning of this VLM for the HPE task fails to achieve desirable HPE accuracy, while some model merging methods can improve accuracy but frequently produce blended invalid response formats, struggling to handle both object detection and HPE tasks simultaneously. To integrate HPE capability into CogVLM effectively, we develop a novel LoRA layer-based model merging method. This merging approach applies a high cosine similarity threshold and a “winner-takes-all” layer selection strategy, aligning attention to the HPE task while preserving original object detection knowledge. It successfully resolves issues with blended invalid response formats and improves accuracy. Results show that our HPE-CogVLM achieves a 31.5% reduction in Mean Absolute Error over the current state-of-the-art CNN model, 6DRepNet, in cross-dataset evaluation. Furthermore, HPE-CogVLM outperforms both directly LoRA fine-tuned and task arithmetic-based merged VLMs across all HPE metrics.
Index Terms:
Vision language model, Model merging, Head pose estimation, Visual grounding task, Catastrophic forgetting problem.
I Introduction
@@
Nowadays, the head pose estimation (HPE) technique is applicable in various fields such as attention estimation [1, 2, 3], face recognition [4, 5, 6, 7], customer behavior analysis [8, 9], driver assistance systems [10, 11, 12, 13] and human-robot interaction [14]. This task involves predicting the Euler angles (yaw, pitch, and roll) of human heads from images or videos. Recent research efforts on some CNN-based models like 6DRepNet [15], HopeNet [16] and WHENet [17] have made significant advancements in HPE.
@@
Despite the recent surge of interest in HPE, the application of this technique still faces several challenges in real-world scenario. The CNN-based models typically rely on narrowly focused datasets such as 300W-LP [18] for training, and validate on similarly constrained datasets like AFLW2000 [18] and BIWI [19]. These datasets primarily feature close-up images focused on a single head, mostly displaying frontal faces with yaw angles from $-99^{\circ}$ to $99^{\circ}$, instead of covering the full range of head poses ranged from $-180^{\circ}$ to $180^{\circ}$. Additionally, the frequent use of close-up images in these datasets reduces input variability, leading to uniform backgrounds. This uniformity could result in limited robustness in diverse real-world environments. The DirectMHP [20] model attempts HPE prediction in a one-shot manner trained on the full-range HPE datasets, like Agora [21] and CMU Panoptic [22], but this model lacks stability and struggles to balance the head bounding box (BBox) detection and HPE task performance. Consequently, the model’s effectiveness remains uncertain in real-world settings.
@@
The appearance of Large Language Models (LLMs) has made substantial advancements in a wide range of applications, significantly enhancing our daily lives by offering sophisticated assistance across various tasks. Recently, Vision Language Models (VLMs) have attracted significant attention for their capability to process multimodal information [23, 24, 25, 26]. By combining image and video understanding with language processing, VLMs can effectively perform complex tasks, such as visual question answering [23, 27, 25] and visual grounding [26, 28]. In this paper, we leverage the attention mechanism of VLMs to enhance the robustness and accuracy of the HPE task. The VLMs empower the model to process the entire image while focusing specifically on the head region, eliminating the need to crop out backgrounds. This capability reduces the risk of overfitting to a limited set of visual features and allows the model to leverage context from the full scene. As a result, it enhances the robustness of tasks that traditional CNN-based methods often struggle to address. We validate this approach by integrating HPE functionality into the grounding model of CogVLM [26]. The grounding CogVLM’s capabilities include caption grounding, referring expression generation, referring expression comprehension and grounded visual question answering [26]. All of these functionalities involve in the description of object localization in the BBox format of [[$x_{0}$, $y_{0}$, $x_{1}$, $y_{1}$]] as shown in Figure 1(a). This BBox grounding capability provides a foundational skill for learning the new HPE task introduced in this paper. By leveraging this capability in designed prompts, our approach can accurately identify the head position within the image, even when multiple people are present in the image.
@@
Despite its advantages, incorporating the HPE task into the grounding CogVLM introduces several challenges. First, VLM tasks such as image description, visual reasoning, and visual perception usually contain answering questions with natural language responses. In contrast, our HPE task requires the VLM to produce precise numerical Euler angles. Although the grounding CogVLM can predict BBoxes, indicating its ability to produce numerical responses, the HPE task is significantly more complicated. HPE requires predicting the human head’s orientation in terms of yaw, pitch, and roll angles, which involves interpreting 3D orientation from 2D images. This introduces additional dimensions of depth and angular perspective not required in the basic BBox detection task. Therefore, it raises the challenge of whether the grounding model can provide HPE answers with much higher accuracy. Secondly, catastrophic forgetting [29, 30, 31] poses a significant challenge in fine-tuning LLMs. The catastrophic forgetting problem is a phenomenon that LLMs tend to forget previously learned information when acquiring new data. Although extensive research has been conducted on mitigating catastrophic forgetting in general tasks, there is currently a lack of research specifically addressing this issue within the context of complex grounding tasks. Lastly, the original grounding CogVLM only involves in outputting responses with natural languages and BBoxes in [[$x_{0}$, $y_{0}$, $x_{1}$, $y_{1}$]] format. In this paper, we introduce a new format {$yaw\_angle$, $pitch\_angle$, $roll\_angle$} for answering HPE prompts as shown in Figure 1(b). This enriches the knowledge of the original grounding CogVLM, meanwhile increasing the complexity of output formats. Empirically, we have observed that the directly LoRA [32] fine-tuning and model merging methods frequently generate blended invalid outputs like [[$x_{0}$, $y_{0}$, $yaw\_angle$}, which is referred as invalid answers in this paper. More invalid answers are detailed in Table II.
@@
Figure 1: Examples of CogVLM and HPE-CogVLM. (a) shows an example of CogVLM grounding capability, which demonstrates the original grounding CogVLM’s ability to identify objects based on prompts, a foundational skill useful for HPE task. (b) displays a visualization of head orientation predicted by our HPE-CogVLM from the CMU Panoptic dataset, using Euler angles. The head pose labels are depicted with pitch (red axis), roll (green axis), and yaw (blue axis) angles, each indicated in their respective directions.
@@
In this paper, for addressing the catastrophic forgetting problem in grounding tasks, we evaluate and improve the data rehearsal methods [29, 30] that were originally used in non-grounding VLMs. The results show that the visual grounding task, which demands accurate numerical outputs, requires a significantly larger rehearsal ratio than non-grounding VLMs. Here, the rehearsal ratio represents the percentage of images randomly selected from earlier training phases that are reintegrated during the training of new tasks [29, 30]. To improve HPE accuracy and address blended invalid outputs, we propose and validate a model merging method based on LoRA layers. Utilizing this approach, our model demonstrates exceptional robustness, achieving a 31.5% reduction in Mean Absolute Error (MAE) of Euler angles, compared to the CNN-based state-of-the-art (SOTA) in cross-dataset evaluations. Furthermore, we compare our LoRA layer-based merged model with both the directly LoRA fine-tuned model and the task arithmetic-based merged model within CogVLM. Our approach consistently shows superior performance in both of MAE and invalid answer ratio reduction. Our contributions can be concluded as following:
@@
  • Our work pioneers the improvement of HPE tasks through leveraging the visual grounding capability of CogVLM, demonstrating the VLM’s ability to manage complex 3D spatial perception while retaining existing object localization knowledge.
  • To the best of our knowledge, this is the first work to explore the issues of catastrophic forgetting and blended invalid response when multiple grounding tasks are involved.
@@
  • We propose a novel LoRA layer-based model merging method that adopts a “winner takes all” strategy, significantly outperforming the CNN-based SOTA, directly LoRA fine-tuned VLM, and task arithmetic-based merged VLM in terms of MAE and invalid answer ratio reduction. This demonstrates our method is able to achieve outstanding robustness and effectiveness in the HPE task, and holds potential for broader application in various grounding tasks.
II Related Work
II-A Head Pose Estimation (HPE)
Traditional approaches for HPE include both landmark-based [33] and landmark-free methods [34, 35]. Landmark-based methods rely on detecting specific facial landmarks, such as the eyes, nose, and mouth, to estimate head orientation. While these methods perform well under controlled conditions, they struggle with full-range HPE because extreme head poses often obscure facial features, making it challenging to detect landmarks accurately [36]. Given our focus on full yaw range HPE, we prioritize landmark-free approaches, which are better suited for this task. Under this approach, several models divide continuous rotation variables into discrete bins for classification purposes [16, 17, 37, 38, 39]. Besides those, FSA-Net [40] employs a stage-wise regression and feature aggregation scheme to predict Euler angles. Meanwhile, models such as 6DRepNet [15] and TriNet [41] take a different approach by estimating the rotation matrix instead of directly predicting Euler angles. Despite their potential, CNN-based methods face significant robustness issues in real-life scenarios, often struggling with lighting variations, occlusions, and complex backgrounds. In this paper, we use several traditional CNN-based landmark-free HPE approaches as baselines. This allows us to assess conventional limitations and benchmark the effectiveness of our proposed method.
II-B Grounding in Vision Language Models
Some VLMs with grounding capabilities can provide accurate BBox information in the format of [[$x_{0}$, $y_{0}$, $x_{1}$, $y_{1}$]] based on specific prompts [26, 42, 43, 44]. This functionality is crucial for tasks requiring precise spatial awareness, like object detection and image captioning, as accurately identifying object positions enhances visual scene understanding. Unlike traditional BBox prediction task, HPE task requires a understanding of 3D spatial relationships to produce accurate Euler angles in the format of {$yaw\_angle$, $pitch\_angle$, $roll\_angle$}. Most existing VLMs are not inherently designed to handle such queries, as they typically output 2D BBoxes rather than 3D rotational data. Currently there is limited research investigating the effectiveness of VLMs in HPE task. While some related efforts exist, such as the work on CLIP-Gaze [45], which focuses on the gaze estimation task by using CLIP [46] model, CLIP-Gaze is limited to applications requiring only a narrow yaw angle range and does not address catastrophic forgetting, leading to a loss of CLIP’s multi-task capabilities.
II-C Model Merging in LLMs
There has been extensive exploration of model merging techniques designed to enhance the capabilities of LLMs. These methods aim to combine multiple LLMs, each with specialized functionalities, into a unified model capable of addressing a range of tasks across various domains. The typical merging methods [47, 48, 49, 50, 51, 52, 53] usually apply rules or algorithms to trim or merge the parameters of LLMs. For example, task arithmetic [47] defines arithmetic rules to incorporate new capabilities or delete undesired ones, allowing task-specific parameters to be adjusted in a controlled manner. Other approaches, like evolutionary model merging [53] leverage evolutionary algorithms to iteratively optimize the merging process. These algorithms evaluate multiple merging configurations across generations, selecting the most effective combinations based on defined fitness criteria. However, we have empirically found that existing merging methods usually produce blended invalid outputs, when dealing with multiple grounding tasks that require distinct output formats. To address this challenge, we propose a novel model merging method designed specifically to mitigate blended invalid outputs.
II-D Catastrophic Forgetting Problem
Catastrophic forgetting has been a significant issue that limits the effectiveness of LLMs, as they tend to forget previously knowledge when learning new knowledge. This problem is particularly pronounced in continual learning settings, where models are sequentially exposed to new tasks and risk losing their ability to perform earlier ones [54]. The rehearsal method [29, 30, 31, 55, 56] is the most widely used method to mitigate catastrophic forgetting. It involves reusing a small portion of old task datasets into the new task fine-tuning process. By periodically revisiting earlier knowledge, the model maintains a balanced representation of all tasks, reducing the likelihood of forgetting. Despite these advancements, catastrophic forgetting remains under-explored in grounding tasks, which require precise numerical outputs. In this paper, we evaluate and refine the rehearsal method for grounding tasks, aiming to balance prior knowledge retention with effective adaptation to new tasks.
III HPE-CogVLM Framework
@@
The proposed framework of HPE-CogVLM as shown in Figure 2 is structured through a multi-stage process. Each stage of this framework is designed to enhance different aspects of the model’s capabilities, gradually refining the parameters to balance HPE and BBox tasks. The fine-tuning process at each stage follows the CogVLM’s fine-tuning scripts [^1] https://github.com/THUDM/CogVLM, which implement LoRA [32] across transformer blocks, including the query, key, value of attention layers and dense layers. Subsequently, the LoRA matrices of each layer are accumulated into the corresponding layer in original model. To support this multi-stage process, the framework utilizes a variety of datasets, each serving a distinct role in model training and evaluation. Table I provides an overview of how these datasets contribute at different stages, such as enhancing human head BBox detection, improving HPE task accuracy, and preventing catastrophic forgetting. Below is a detailed description of each stage in the framework:
TABLE I: A detailed overview of various datasets used in our framework.
Task Dataset # of Images # of Heads Usage Functionality
Train Test Train Test
Weak Label Images CrowdHuman [57] 11,731 - 94,795 - Stage 1
Focus on human head BBox detection
Warm-up for HPE task using weak label images
Task-specific Images Agora [21] 9,654 - 64,187 - Stage 2, 4
Training with precise HPE labels from synthetic images
Focus on HPE task to improve accuracy
Rehearsal Images Refcoco [28] 42,404 - - - Stage 4 Preventing catastrophic forgetting with Refcoco/+/g training set Investigating rehearsal ratio optimization
Refcoco+ [28] 42,278 - - -
Refcocog [58] 42,226 - - -
Evaluation Images CMU Panoptic [22] - 16,216 - 32,738 Stage 5 Evaluating the accuracy of the HPE task on CMU Panoptic test set Evaluating the BBox detection knowledge on Refcoco/+/g test set
Refcoco [28] - 3785 - -
Refcoco+ [28] - 3773 - -
Refcocog [58] - 5023 - -
@@
III-A Stage 1: Pre-training of the Original Grounding CogVLM on Weak Label Data
As indicated in the CrowdHuman dataset [57] functionality column of Table I, the primary goal at this initial stage is to train the model to develop its capability for human head BBox detection. Additionally, this stage acts as a warm-up for the model’s HPE capability by utilizing weak label images. This prepares the model to initiate the HPE task in real-world scenarios.